Given r1 = 3i − 4j + 3k, r2 = 5i + 3j − 6k, r3 = 2i + 7j + 3k and r4 = 4i + 3j + 5k.





Find the scalars a, b and c such that r4 = ar1 + br2 + cr3







1
Expert's answer
2022-06-03T12:34:14-0400
"3a+5b+2c=4"

"-4a+3b+7c=3"

"3a-6b+3c=5"

"A=\begin{pmatrix}
 3 & 5 & 2 & & 4 \\
 -4 & 3 & 7 & & 3 \\
 3 & -6 & 3 & & 5 \\
\end{pmatrix}"

"R_1=R_1\/3"


"\begin{pmatrix}
 1 & 5/3 & 2/3 & & 4/3 \\
 -4 & 3 & 7 & & 3 \\
 3 & -6 & 3 & & 5 \\
\end{pmatrix}"

"R_2=R_2+4R_1"


"\begin{pmatrix}
 1 & 5/3 & 2/3 & & 4/3 \\
 0 & 29/3 & 29/3 & & 25/3 \\
 3 & -6 & 3 & & 5 \\
\end{pmatrix}"

"R_3=R_3-3R_1"


"\begin{pmatrix}
 1 & 5/3 & 2/3 & & 4/3 \\
 0 & 29/3 & 29/3 & & 25/3 \\
 0 & -11 & 1 & & 1 \\
\end{pmatrix}"

"R_2=3R_2\/29"


"\begin{pmatrix}
 1 & 5/3 & 2/3 & & 4/3 \\
 0 & 1 & 1 & & 25/29 \\
 0 & -11 & 1 & & 1 \\
\end{pmatrix}"

"R_1=R_1-5R_2\/3"


"\begin{pmatrix}
 1 & 0 & -1 & & -3/29 \\
 0 & 1 & 1 & & 25/29 \\
 0 & -11 & 1 & & 1 \\
\end{pmatrix}"

"R_3=R_3+11R_2"


"\begin{pmatrix}
 1 & 0 & -1 & & -3/29 \\
 0 & 1 & 1 & & 25/29 \\
 0 & 0 & 12 & & 304/29 \\
\end{pmatrix}"

"R_3=R_3\/12"


"\begin{pmatrix}
 1 & 0 & -1 & & -3/29 \\
 0 & 1 & 1 & & 25/29 \\
 0 & 0 & 1 & & 76/87 \\
\end{pmatrix}"

"R_1=R_1+R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 67/87 \\
 0 & 1 & 1 & & 25/29 \\
 0 & 0 & 1 & & 76/87 \\
\end{pmatrix}"

"R_2=R_2-R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 67/87 \\
 0 & 1 & 0 & & -1/87 \\
 0 & 0 & 1 & & 76/87 \\
\end{pmatrix}"


"a=\dfrac{67}{87}, b=-\dfrac{1}{87}, c=\dfrac{76}{87}"